OpenStudy (anonymous):

Find an equation for the parabola that has its vertex at the origin and satisfies the given condition. Focus F 0, − 1 2

OpenStudy (dumbcow):

Focus at (0,-12) ?? if so then -12 = 1/4a where y = a(x)^2

OpenStudy (anonymous):

0, -1/2 sorry about that

OpenStudy (dumbcow):

ok -1/2 = 1/(4a) where y=ax^2